A particle is projected with velocity ų making an angle of 45°wiyh vertical. to what height does it rise?

Grade:11

1 Answers

Nishant Vora IIT Patna
7 years ago
MAX height of projectile is u^2 sin^2 @ / 2g = u^2 (1/2) / 2 g = u^2 / 4g